Cepeda Manufacturing Company is considering three new projects, each requiring an equipment investment of $22,000. Each project will last for 3 years and produce the following cash inflows. Year AA BB CC 1 $7,000 $9,500 11,000 9,000 9,500 10,000 15,000 9,500 9,000 Total $31,000 28,500 $30,000 The equipments salvage value is zero. Cepeda uses straight-line depreciation. Cepeda will not accept any project with a payback period over 2 years. Cepedas minimum required rate of return is 12%.
Answer #1

CALCULATE PAYBACK PERIOD :

PAYBACK PERIOD = INTIAL INVESTMENT/ANNUAL CASH FLOW

PROJECT AA          = 2 YEAR+6000/15000   = 2.4 YEARS

PROJECT BB          = 22000/9500 = 2.32 YEARS

PROJECT CC           = 2 YEAR+1000/9000 = 2.11 YEARS

MOST DESIREABLE PROJECT IS PROJECT CC

LEASE DESIRABLE PROJECT IS PROJECT AA

BUT ACCORDING TO COMPANY'S POLICY THERE SHOULD NOT ACCEPTED ANY PROJECT BECAUSE ALL PROJECT'S PAYBACK IS OVER 2 YEARS